OpenStudy (anonymous):

Karrie bought a refrigerator from a rental center for $989.47. She makes 12 monthly payments of $97.32 with her credit card. The rental center charges $2.25 for every purchase made with a credit card. She also had to pay late fees in the amount of $55 four different times. What is the total cost of the refrigerator?

OpenStudy (anonymous):

well you know that the renatl centers charges will be 2.25 multiplied by 12 since she is making twelve payments

OpenStudy (anonymous):

is it 1414.84

OpenStudy (agent0smith):

@Hope_nicole it only charges 2.25 for a purchase.

OpenStudy (anonymous):

ohh my bad, thank you @agent0smith .

OpenStudy (agent0smith):

12 payments of 97.32, 4 late fees of 55, and one 2.25 purchase fee: \[12 \times $97.2 + 4 \times $55 + 1 \times $2.25 = ?\]

OpenStudy (anonymous):

make sure to do your multiplication first @beckyabney

OpenStudy (agent0smith):

is it 1414.84... close but not quite.

OpenStudy (anonymous):

@agent0smith that would mean the answer is 1390.09 but thats not even an answer...

OpenStudy (agent0smith):

Closer... but still not quite correct. \[\large (12 \times $97.2) + (4 \times $55) +( 1 \times $2.25) = ?\]

OpenStudy (anonymous):

its 97.32 not 97.2. i typed that same exact equation into my calculater and got 1390.09 and thats not an option

OpenStudy (agent0smith):

Oh, my bad, i misread it. What are the answer choices?

OpenStudy (anonymous):

$1,387.84 $1,414.84 $1,167.84 $1,194.84

OpenStudy (agent0smith):

Well then it looks like the $2.25 doesn't count as part of the cost, for whatever reason. If you remove the 2.25, you get one of those answers.

OpenStudy (agent0smith):

Or if you remove the late fees as well you get another of those answers :/

OpenStudy (anonymous):

i think that they charge 2.25 for every credit card charge

OpenStudy (agent0smith):

If the credit card actually charges 2.25 for every *payment* then the question is just very poorly worded.

OpenStudy (agent0smith):

The question states "$2.25 for every purchase made with a credit card" the only purchase she made is the fridge.